I would be so grateful if you could assist. I have a very simple question, I believe, which is as follows.
A Hebrew word consists of 3 consonants, and the Hebrew alphabet of 22 consonants.
How many permutations of 3 can be extracted from 22?
Also how many permutations of 2 consonants and how many of 4?
The answer should provide a theoretical maximum for the number of Hebrew words in existence, now and/or in the past. The numbers will form part of some research on the Hebrew origins of English.
I hope you can help. I have done a calculation myself, but don't trust my own judgement, I regret. Such was algebra when I was at school.
Danny Israel.



Found 2 solutions by Edwin McCravy, richwmiller:
Answer by Edwin McCravy(20054)

A Hebrew word consists of 3 consonants, and the Hebrew alphabet of 22 consonants.
How many permutations of 3 can be extracted from 22?
I will assume the consonants do not have to be different.
That is, they could be like XXZ, XZZ or even XXX.

Then there would be 26×26×26 = 17,576

Also how many permutations of 2 consonants
22×22 = 484

and how many of 4?
22×22×22×22 = 234,256

The numbers would be fewer if there are rules about using the 
same letter.  FI, if, say, as in English, there cannot be more 
than 2 of the same letter in a row.  Also in English, words do
not end in j, v, or q, and few end in vowels other than e. Few
words begin with a double letter. q must be followed by u.
Many spelling rules will reduce the numbers.  Also, are there 
no vowels in Hebrew? 

Edwin

Answer by richwmiller(17219)

I believe vowels can be written or not (as you wish) in Hebrew or Arabic.